The length of a rectangular garden is 4 feet more than three times its width. The perimeter of the garden is 56 feet. Find the length and width of the garden.
Answer by hamsanash1981@gmail.com(151)

Let Width be x and length 3x+4
Perimeter = 56
2(l+b)=56
2(x+3x+4)=56
2(4x+4)=56
4x +4 = 28
4x = 24
x = 6
Width = 6 ft and Length = 22 ft